# Taming Data and Transformers for Audio Generation

### v2

@@ 4.2 Text-2-Audio Generation
Model Realism Quality Prompt Alignment Overall Preference
GenAU-L vs GenAU-S 61.20% 58.00% 61.20% 60.40%
GenAU-L vs GenAU-L (AC) 60.40% 54.80% 60.40% 59.20%
GenAU-L vs MAD-2 64.00% 62.40% 68.40% 66.40%
GenAU-S w/o Recap. vs MAD-2 64.40% 64.-0% 63.20% 64.80%
Metrics. We compare the performance of our method with baselines using the standard Frechet Distance (FD), Inception score (IS), and CLAP score on the Audioset test dataset, containing 964 samples. There is little consistency between baselines when computing the metrics. Some prior work reported the Fréchet distance results using the VGGish network (Hershey et al., 2017), denoted as (FAD) (Kilgour et al., 2019), while other uses PANNs (Kong et al., 2019). Additionally, to compute the CLAP score, some prior work (Liu et al., 2023c) used CLAP from LAION, which we denote as $\text{CLAP}_{LAION}$ (Wu et al., 2023), while others (Majumder et al., 2024; Huang et al., 2023b; a) used CLAP from Microsoft (Elizalde et al., 2023), which we denote as $\text{CLAP}_{MS}$. Furthermore, some prior (Liu et al., 2023b; c) used CLAP re-ranking with 3 samples for computing the metrics. Due to such inconsistencies in evaluation pipelines and varying results for the same baselines reported in different studies, we recompute all metrics using the official checkpoints to ensure consistent comparisons. We follow the same evaluation protocols of AudioLDM (Liu et al., 2023b) without CLAP re-ranking and use the AudioLDM evaluation package to compute the metrics. Besides, we run our ablations on the Bigsoundbank split from WavText5k (Deshmukh et al., 2022), which serves as an out-of-distribution evaluation for our models. This is to prevent biasing the evaluation based on the training data. Finally, to further validate our results we run a user preference study. Details about the user study can be found in the Appendix.
Results. In Tab. 4, we report evaluation results. Our method achieves superior performance compared to the state-of-the-art methods in terms of IS, FAD, $\text{CLAP}_{MS}$ and $\text{CLAP}_{LAION}$ scores, marking an improvement of $22.7\%$, $4.7\%$, $3.6\%$, and $13.5\%$, respectively. This shows that GenAu can produce high audio quality and achieve better semantic alignment with the conditioning text.
Data scaling. We consider two key aspects: data quality and quantity. First, in Tab. 5 ($2^{nd}$ fow), we show that GenAu-L trained with AutoReCap is generally favoured over training only with AudioCaps (AC). This is confirmed in Tab. 3 ($1^{st}$ vs $2^{nd}$ row), where increasing the dataset size significantly boosts the results across all metrics, improving IS by $56.3\%$. Additionally, we show ($3^{rd}$ vs $4^{th}$ row) that using AutoCap to recaption the dataset significantly enhances the results over all metrics, confirming the importance of data quality. Interestingly, expanding the data size at a lower caption quality does not yield similar gains even at a bigger model ($2^{nd}$ vs $4^{th}$ row), aligning with results reported by Liu et al. (2023c). This highlights that data quality brought by AutoCap is as crucial as the data quantity. Lastly, we examine the effect of scaling the data with synthetic captions. For this, we train for 50k steps by fixing AC and Clotho in the training data and varying the amount of synthetic data. As reported in Fig. 3 (right), increasing the amount of data with synthetic caption has a clear improvement over both IS and FD, with the model trained on the whole AutoReCap achieving the best results.
Model size scaling. In Tab. 5, we report ($1^{st}$ row) that GenAu-L (1.25B params) is constantly favoured over GenAu-S (493M params). This is further confirmed by our automatic evaluation in Tab. 3 ($1^{st}$ vs $3^{rd}$ row), where the larger model shows significant improvements across all metrics. The scaling trend is also evident in Fig. 3, which demonstrates a clear correlation between model size and performance in terms of both IS and FD scores.
Model architecture ablation. Until recently, A UNet (Ronneberger et al., 2015) has been the most popular choice for the diffusion backbone. Yet, as reported in Tab. 4, replacing the FiT backbone with a UNet drastically reduces performance across all metrics. This supports baseline findings where UNet-based methods lag behind transformer-based approaches (Huang et al., 2023a). Another choice that has recently gained popularity is the DiT architecture (Peebles & Xie, 2023). Make-an-Audio-2 (MAD-2) employs a DiT at a similar model size and data scale as GenAU-L. However, as we show in Tab. 5, our model is consistently preferred over MAD-2 ($3^{rd}$ row), even without dataset recaptioning ($4^{th}$ row) (i.e. at similar data settings). We infer that the FiT architect, with its read and write operations, allocates compute more efficiently to the key segments of the input, making it more suitable to ambient audio clips which often include silent or redundant parts.
@@ 5 Conclusion
We take a holistic approach to improve the quality of existing audio generators. Starting by addressing the scarcity of large-scale captioned audio datasets, we build a state-of-the-art audio captioning method, AutoCap, which leverages audio metadata to collect a dataset of 57M annotated audio clips. We then built a latent diffusion model based on a scalable transformer architecture which we trained on our re-captioned dataset to obtain GenAu, a state-of-the-art open-sources model for audio generation. Our approach not only improves ambient audio generation but also opens up possibilities for extending GenAu to other domains, such as speech and music generation. As an additional contribution, we built AutoReCap-XL, a text-audio-video ambient audio dataset with an unprecedented size of $57M$ pairs. AutoReCap-XL can potentially serve as a joint text-audio-video dataset and broadens novel applications such as text-to-audio-video joint generation, a more natural and desired choice of video generation. We leave larger analyses on this dataset for future work.
@@ A.2 Stage 2: Speech and Music Filtering
We downloaded English transcripts from YouTube and used automatically generated ones for videos without existing transcripts. However, we discard videos without any transcripts. While some datasets provide only video segments with specific timestamps, we processed the full videos, totaling around 73 million videos. We accepted audio segments longer than one second that lacked any corresponding subtitles, indicating the absence of speech and music. After filtering, we isolated approximately 327.3 million segments from 55.1 million videos. Fig. 5 displays the distribution of the number of segments per video. We denote this dataset as AutoReCap-XL-Raw. Subsequently, we use AutoCap to caption the audio segments. Fig. 7 shows the distribution of caption lengths. Given that AutoReCap was trained for 10-second audio, we limited segments to this duration. Additionally, we concatenate consecutive segments yielding identical captions to form longer audio clips. Fig. 9 illustrates the audio length distribution, and a word cloud of the captions is shown in Fig. 11. Despite filtering, the dataset was still dominated by captions related to speech and music. We attribute this to the limitations of YouTube’s automatic transcription, particularly with certain types of music and less common languages.
@@
To further refine the dataset from speech and music, We follow a simple filtering approach. Specifically, we employed a large language model (LLM) to generate keywords associated with speech and music, such as "talking", "speaking", and "singing," and excluded all audio segments whose captions contained such keywords. This process yielded 57 million audio-text pairs from 20.3 million videos. Fig. 5 shows the number of segments per video, Fig. 7 shows the caption length distribution, Fig. 9 shows the audio length distribution, and Fig. 11 presents a word cloud of the final captions. We outline the data sources for constructing this dataset in Tab. 6. Our proposed dataset is not only 90 times larger than the previously largest available dataset, LAION-Audio-630KWu et al. (2023) in terms of the number of audio clips, but also provides more accurate captions compared to existing datasets that rely on raw textual data. A comprehensive comparison with other datasets is detailed in Tab. 7
Appendix B Limitations.
B.1 AutoCap
@@ B.3 AutoReCap-XL
Appendix C Evaluation Details
C.1 Audio Captioning
While the established practice in the evaluation of audio captioning methods is to report the results on the test set using the checkpoint that performs best on the validation subset, prior work (Étienne Labbé et al., 2023; Kim et al., 2024b) reported high instability of the metrics on the validation subset and weak correlation between the validation and test performance, making the model’s results vary significantly for different seeds. To alleviate this, ENCLAP (Kim et al., 2024b) selects around five best-performing validation checkpoints and reports their best results on the test set. CNext-trans (Étienne Labbé et al., 2023) uses the FENSE score to pick the best validation checkpoint. This method of choosing the best checkpoint may produce misleading results and potentially disadvantage baselines. Our model, thanks to the two-stage training paradigm, significantly reduces this instability and we observe steady performance gains as training progresses. Therefore, we report the results at convergence, specifically after 20 epochs of pre-training and 20 epochs of fine-tuning.
C.2 Audio Generation
There is a lack of consistency in the metrics used across text-to-audio generation baselines. Some baselines, such as Liu et al. (2023b) and Huang et al. (2023a), employ the VGGish network (Hershey et al., 2017) to compute the Fréchet Distance, while others, like Liu et al. (2023c), utilize the PANNs network (Kong et al., 2019), and still others rely on OpenL3 embeddings, such as Evans et al. (2024b). Additionally, some baselines use the LAION CLAP network (Wu et al., 2023) to compute the CLAP score, whereas others use the Microsoft CLAP network (Elizalde et al., 2023). To further complicate matters, different baselines often report varying results in various publications. To address these discrepancies, we recalculated all metrics for the baselines using their publicly released checkpoints under identical evaluation configurations. Our method significantly outperforms the baselines across all metrics, except for the Fréchet Distance, where it is slightly behind Make-An-Audio 2 (Huang et al., 2023a). Nevertheless, our user study, detailed in the main paper, indicates that GenAu is generally preferred over Make-An-Audio 2.
@@ C.2 Audio Generation
AutoReCap-XL 57M 123,500 Automatic re-captioning
@@
Prompt Alignment Considering the prompt to generate the audio is "A sewing machine operating as a machine motor hisses loudly in the background", which audio better follows the given prompt?
Overall Preference Considering the realism, quality, and prompt alignment of the audio, which audio do you prefer more overall? The prompt is: "A sewing machine operating as a machine motor hisses loudly in the background."
C.3 User Study
@@
Appendix D Training and Inference Details
D.1 AutoCap

### v3

@@ 4.1 Automatic Audio Captioning
GenAu-Large-Full 1.25B 19.6M 14.48 12.56 1.71 0.318 0.677
Results. Tab. 1 reports that our method outperforms baselines on all metrics, achieving notable improvements in CIDEr (83.2) and BLUE1 (73.1) scores. Notably, even without metadata (i.eaudio only), AutoCap surpasses baselines in most metrics. We found that incorporating metadata significantly enhances CIDEr but slightly reduces SPICE. This trade-off likely results from the enhanced descriptive detail brought by the metadata, which while enriching the content, introduces noise that may compromise the model’s semantic precision. In addition, AudioCaps is labeled based on audio alone. Thus, the evaluation penalizes the description of information that can not be deduced with certainty from the audio modality only, such as the specific type of object producing the sound. Qualitatively, our captions are more detailed and temporally accurate than baselines. ENCLAP-Large often misses key details. CNext-trans, while accurate, often produces short captions that lack details. We include qualitative comparisons in the Website and Appendix. Moreover, AutoCap is four times faster than ENCALP, producing a caption for a 10-second clip in 0.28 seconds, compared to ENCALP which takes 1.12 seconds. Furthermore, we observe consistent improvements when pretraining on weakly-labeled data, validating the effectiveness of our training strategy in benefiting from larger, weakly-labeled datasets.
Ablations. In Tab. 2, we ablate model design choices. Using CLAP embedding brings a $2.5$ points increase in CIDEr. Omitting Stage 2 training, which involves finetuning BART (Lewis et al., 2020), results in performance degradation, likely due to the necessity of adapting BART’s decoder to the sentence structure typical of AudioCaps. A more severe degradation in performance is observed when Stage 1 is not performed, as the misaligned representation between the encoder and the decoder leads to catastrophic forgetting in the language model. Finally, finetunning BART word embeddings in Stage 2 reduces performance slightly.
4.2 Text-2-Audio Generation
Training dataset and details. We follow baselines (Liu et al., 2023b; Huang et al., 2023a; Majumder et al., 2024) and train on 10-second clips at 16kHz resolution. We use a patch size of 1 and a group size of 32. We use LAMB optimizer (You et al., 2019) with a LR of 5e-3. We train for 220k steps and choose the checkpoint with the highest IS.
Baselines. We compare with TANGO 1 & 2, (Ghosal et al., 2023), AudioLDM 1 & 2 (Liu et al., 2023a, b), and Make-An-Audio 1 & 2 (Huang et al., 2023b, a). Both AudioLDM and Make-an-Audio train a UNet-based latent diffusion model (Rombach et al., 2022) on Mel-Spectrogram representation, by regarding it as a single channel image, and use a pretrained CLAP encoder to condition the generation on an input prompt. TANGO proposed to use FLAN-T5 (Chung et al., 2024) as the text encoder and reported significant improvements. AudioLDM-2 and Make-an-Audio-2 proposed to use a dual encoder strategy of a T5 (Raffel et al., 2022) and CLAP encoder. Make-an-Audio-2 proposes to use a 1D VAE representation and employ a DiT as the diffusion backbone. Recently, Tango-2 proposed to use instruction fine-tuning on a synthetic dataset to enhance temporal understanding. In our experiments, we focus on text-conditioned natural audio generation.
Metrics. We compare the performance of our method with baselines using the standard Frechet Distance (FD), Inception score (IS), and CLAP score on the AudioSet test split. There is little consistency between baselines when computing the metrics. Some prior work reported the Fréchet distance results using the VGGish network (Hershey et al., 2017), denoted as (FAD) (Kilgour et al., 2018), while other uses PANNs (Kong et al., 2019). Additionally, to compute the CLAP score, some prior work (Liu et al., 2023b) used CLAP from LAION, which we denote as $\text{CLAP}_{LAION}$ (Wu et al., 2023b), while others (Majumder et al., 2024; Huang et al., 2023b, a) used CLAP from Microsoft (Elizalde et al., 2023), which we denote as $\text{CLAP}_{MS}$. Furthermore, some prior (Liu et al., 2023a, b) used CLAP re-ranking with 3 samples for computing the metrics. Due to such inconsistencies in evaluation pipelines and varying results for the same baselines reported in different studies, we recompute all metrics using the official checkpoints to ensure consistent comparisons. We follow the same evaluation protocols of AudioLDM (Liu et al., 2023a) without CLAP re-ranking and use the AudioLDM evaluation package to compute the metrics. Besides, to prevent biasing the evaluation based on the training data, we run our ablations on the Bigsoundbank split from WavText5k (Deshmukh et al., 2022), which serves as an out-of-distribution evaluation for our models. Finally, to further validate our results, we run user studies. Details about the user study can be found in the Appendix.
Table 4: GenAu ablation study on out-of-distribution dataset.
Ablation Model IS FD $\text{CLAP}_{MS}$
Model Scale
GenAU-S 15.76 21.29 0.36
GenAU-L 18.98 20.81 0.38
Data Scale
GenAU-L (AC) 12.14 25.82 0.30
GenAU-L (AutoReCap) 18.98 20.81 0.38
Synthetic Captions
GenAU-S w/o Recap. 11.83 25.34 0.29
GenAU-S 15.76 21.29 0.36
Backbone
GenAU-S w/ U-Net 13.55 32.51 0.28
GenAU-S w/ DiT 13.96 23.32 0.33
GenAU-S w/ FiT 15.76 21.29 0.36
Comparison with baselines. In Tab. 3, we report evaluation results. When trained with a similar size and data scale to state-of-the-art methods, GenAu achieves superior performance in most metrics, improving IS by $11.1\%$, FAD by $4.7\%$, and $13.5\%$ in $CLAP_{LAION}$. Using the full 10-seconds subset of AutoReCap-XL further enhances the results with over 21.8% in IS and 20.5% in CLAP score. Additionally, to isolate the impact of model architecture from data quality, we conduct a user study in Tab. 5, where GenAu is consistently preferred over Make-an-Audio-2, even when trained with comparable data quality.
How GenAu scale with synthetic data? To study this, we train GenAu-S (493M params) for 50k steps by fixing AudioCaps and Clotho in the training data and varying the amount of synthetic data from AutoReCap. As reported in Fig. 3 (right), increasing synthetic training data consistently improves both IS and FD. Similar improvements are evident in Tab. 4 where increasing the dataset size significantly boosts all metrics, improving IS by $56.3\%$. User studies in Tab. 5 further support these findings, where a model training with AutoReCap is consistently favored over a model trained only on AudioCaps. Finally, we report in Tab. 3 the scaling performance on AutoReCap-XL (47M). To ensure a fair comparison, we train on 10-second subset (19.7M) and observe improvements of 12.3% in FD and 11.6% in CLAP score compared to training with AutoReCap (800k).
Is caption quality important? We compare in Tab. 4 GenAu-S against the same model trainined with WavCaps (Mei et al., 2024a) captions. We observe gains across all metrics, confirming the importance of caption quality and the improvements brought by our high-quality captioner AutoCap. Interestingly, expanding data size with lower-quality captions offers no significant gains over training on AudioCaps alone, consistent with Liu et al. (2023b).
Does GenAu benefit from model size scaling? Similar to data scaling, increasing model size consistently enhances performance. As shown in  Fig. 3 (left), larger models achieve better FD and IS. This is further confirmed in Tab. 4, where GenAu-L (1.25B) outperforms GenAu-S (493M) across all metrics with almost 20.5% increase in IS. User study in Tab. 5 shows a strong preference for GenAu-L over GenAu-S. Unlike previous methods which reported diminishing returns with model scaling (Hai et al., 2024; Liu et al., 2023b), GenAu continues to improve as model size scale.
How do FITs compare with other diffusion backbones? We evaluate the impact of the diffusion backbone by replacing FIT with a UNet (Ronneberger et al., 2015), or a DiT (Peebles & Xie, 2023). In Tab. 4, we observe that GenAu with FIT outperforms these alternatives across all metrics. We infer that the FIT architect, with its read and write operations, allocates compute more efficiently to the key segments of the input, making it suitable for ambient audio clips which often include silent or redundant parts.
@@ 4.2 Text-2-Audio Generation
Comparison Realism Quality P-Align. Pref.
Model: GenAU-L vs GenAU-S 61.20% 58.00% 61.20% 60.40%
Data: GenAU-L vs GenAU-L (AC) 60.40% 54.80% 60.40% 59.20%
Backbone: GenAU-S w/o Recap. vs MAD-2 64.40% 64.00% 63.20% 64.80%
@@ 5 Conclusion
We take a holistic approach to improve the quality of existing audio generators. Starting by addressing the scarcity of large-scale captioned audio datasets, we propose a scalable and efficient dataset collection pipeline. We then build a state-of-the-art audio captioning method, AutoCap, which leverages audio metadata to annotate a dataset of 47M annotated audio clips tailored for large-scale audio generation. We then built a latent diffusion model based on a scalable transformer architecture which we trained on our re-captioned dataset to obtain GenAu, a high-quality open-source model for audio generation. Our approach opens up possibilities for extending GenAu to other domains, such as speech and music. Additionally, AutoReCap-XL can serve as a joint text-audio-video dataset and broadens novel applications such as text to audio-video joint generation.
Limitations and future work. AutoCap was fine-tuned on AudioCaps, featuring 4,892 unique words, which limits the diversity of our generated captions. Consequently, GenAu may face challenges in accurately generating audio for detailed prompts. Additionally, while AutoReCap is extensive in size, it has only been validated for audio generation. We leave broader analysis on more tasks for future work.
@@ A.2 Stage 2: Speech and Music Filtering
We downloaded English transcripts from YouTube and used automatically generated ones for videos without existing transcripts. However, we discard videos without any transcripts. While some datasets provide only video segments with specific timestamps, we processed the full videos, totaling around 73 million videos. We accepted audio segments longer than one second that lacked any corresponding subtitles, indicating the absence of speech and music. After filtering, we isolated approximately 327.3 million segments from 55.1 million videos. Fig. 5 displays the distribution of the number of segments per video. We denote this dataset as AutoReCap-XL-Raw. Subsequently, we use AutoCap to caption the audio segments. Fig. 7 shows the distribution of caption lengths. Given that AutoCap was trained for 10-second audio, we limited segments to this duration. Additionally, we concatenate consecutive segments yielding identical captions to form longer audio clips. Fig. 9 illustrates the audio length distribution, and a word cloud of the captions is shown in Fig. 11. Despite filtering, the dataset was still dominated by captions related to speech and music. We attribute this to the limitations of YouTube’s automatic transcription, particularly with certain types of music and less common languages.
@@
To further refine the dataset from speech and music, We follow a simple filtering approach. Specifically, we employed a large language model (LLM) to generate keywords associated with speech and music, such as ”talking”, ”speaking”, and ”singing,” and excluded all audio segments whose captions contained such keywords. Finally, we filter all audio-text pairs with CLAP score less than 0.1. A distribution of the CLAP scores for the collected dataset is present in Fig. 12. This process yielded 47 million audio-text pairs from 20.3 million videos. Fig. 5 shows the number of segments per video, Fig. 7 shows the caption length distribution, Fig. 9 shows the audio length distribution, and Fig. 11 presents a word cloud of the final captions. We outline the data sources for constructing this dataset in Tab. 6. Our proposed dataset is not only 75 times larger than the previously largest available dataset, LAION-Audio-630K(Wu et al., 2023b) in terms of the number of audio clips, but also provides more accurate captions compared to existing datasets that rely on raw textual data. A comprehensive comparison with other datasets is detailed in Tab. 7
Appendix B Architecture details
B.1 HTSAT Embeddings Extraction
AutoCap uses HTSAT (Chen et al., 2022) embeddings to encode the input audio and follows the HTSAT-BART (Mei et al., 2024a) embedding extraction procedure, described in the following, to obtain “fine-grained” HTSAT embeddings. Given a 10-seconds single-channel input audio at 32Khz, HTSAT represents it as a mel-spectrogram using window size of 1024, 320 hop size, and 64 mel-bins, resulting in an input of shape $(T=1024,F=64)$. The spectrogram is then encoded as latent tokens of shape ($\frac{T}{8P}=32$, $\frac{F}{8P}=2$, $7688D=$ 768) before the classification layer. HTSAT-BART (Mei et al., 2024a), then averages over the frequency dimension to obtain a representation of shape ($\frac{T}{8P}=32$, $1$, $7688D=$ 768) and replicates the latent token by a token replication factor of $328P=$ 32 to obtain a so-called “fine-grained” representation of shape $32\times 32\times 768$, which is flattened into a representation of shape $1024\times 768$. We adopt this representation throughout our work, and Appx. G.3 provides additional evaluation results showing the performance benefits of the token replication operation.
Appendix C Limitations
C.1 AutoCap
@@ C.3 AutoReCap-XL
Appendix D Evaluation Details
D.1 Audio Captioning
While the established practice in the evaluation of audio captioning methods is to report the results on the test set using the checkpoint that performs best on the validation subset, prior work (Labb et al., 2024; Kim et al., 2024b) reported high instability of the metrics on the validation subset and weak correlation between the validation and test performance, making the model’s results vary significantly for different seeds. To alleviate this, ENCLAP (Kim et al., 2024b) selects around five best-performing validation checkpoints and reports their best results on the test set. CNext-trans (Labb et al., 2024) uses the FENSE score to pick the best validation checkpoint. This method of choosing the best checkpoint may produce misleading results and potentially disadvantage baselines. Our model, thanks to the two-stage training paradigm, significantly reduces this instability and we observe steady performance gains as training progresses. Therefore, we report the results at convergence, specifically after 20 epochs of pre-training and 20 epochs of fine-tuning.
D.2 Audio Generation
There is a lack of consistency in the metrics used across text-to-audio generation baselines. Some baselines, such as Liu et al. (2023a) and Huang et al. (2023a), employ the VGGish network (Hershey et al., 2017) to compute the Fréchet Distance, while others, like Liu et al. (2023b), utilize the PANNs network (Kong et al., 2019), and still others rely on OpenL3 embeddings, such as Evans et al. (2024a). Additionally, some baselines use the LAION CLAP network (Wu et al., 2023b) to compute the CLAP score, whereas others use the Microsoft CLAP network (Elizalde et al., 2023). To further complicate matters, different baselines often report varying results in various publications. To address these discrepancies, we recalculated all metrics for the baselines using their publicly released checkpoints under identical evaluation configurations. Our method significantly outperforms the baselines across all metrics, except for the Fréchet Distance, where it is slightly behind Make-An-Audio 2 (Huang et al., 2023a). Nevertheless, our user study, detailed in the main paper, indicates that GenAu is generally preferred over Make-An-Audio 2.
@@ D.2 Audio Generation
AutoReCap-XL 47M 123,500 Automatic re-captioning
@@
Prompt Alignment Considering the prompt to generate the audio is ”A sewing machine operating as a machine motor hisses loudly in the background”, which audio better follows the given prompt?
Overall Preference Considering the realism, quality, and prompt alignment of the audio, which audio do you prefer more overall? The prompt is: ”A sewing machine operating as a machine motor hisses loudly in the background.”
D.3 User Study
@@
Appendix E Training and Inference Details
E.1 AutoCap
AutoCap introduces 6.2 million new parameters on top of the frozen HTSAT audio encoder and the base BART model. These parameters include 4.7M for the Q-Former, 0.9M for embedding layers, and 0.6M for projection layers. The Q-Former employs 256 learnable tokens, a hidden dimension of 256, 8 attention heads, and 2 hidden layers.